OpenStudy (anonymous):

if F(x)=3x-9 and g(x)=-2x+2, find ( f o g )(fx)

OpenStudy (anonymous):

It is basically: f(g(f(x)) right>>

OpenStudy (lgbasallote):

(f o g)(f(x) means f(g(f(x) ) ) so first find g(f(x) ) g(f(x) ) = -2 (f(x ) ) +2 now sub in the value of f(x) -2(3x - 9) +2 -6x + 18 +2 -6x + 20 now sub this in into f(x) f(g(f(x) ) ) = 3(-6x + 20) - 9 -18x + 60 - 9 -18x +51 do you get that?

OpenStudy (anonymous):

No. I had a typo. It really is (f o g)(x)

OpenStudy (anonymous):

Then it is f(g(x)).. put g(x) in place of x in f(x) That is: replace x by -2x + 2 in f(x).. f(x) = 3x - 9 f(-2x + 2) = 3(-2x + 2) - 9 f(g(x)) = -6x + 6 - 9 Or, \[(f \;o \; g)(x) = -6x - 3\]

OpenStudy (anonymous):

Thank you !!!!!!!!!!!!!

OpenStudy (anonymous):

Welcome dear..